# Planning in Branch-and-Bound: Model-Based Reinforcement Learning for Exact Combinatorial Optimization

### v1

@@ 2 Problem statement
with $n$ the number of variables, $m$ the number of linear constraints, $l,u\in\mathbb{R}^{n}$ the lower and upper bound vectors, $A\in\mathbb{R}^{m\times n}$ the constraint matrix, $b\in\mathbb{R}^{m}$ the right-hand side vector, $c\in\mathbb{R}^{n}$ the objective function, and $\mathcal{I}$ the indices of integer variables. Throughout this document, we are interested in repeated MILPs of fixed dimension $\{P_{i}=(A_{i},b_{i},c_{i},l_{i},u_{i})\}_{i\in\mathcal{D}}$ sampled according to an unknown distribution $p_{0}:\Omega\rightarrow\mathbb{R}^{m\times n}\times\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$.
In order to solve MILPs efficiently, the B&B algorithm iteratively builds a binary tree $(\mathcal{V},\mathcal{E})$ where each node corresponds to a MILP, starting from the root node $v_{0}\in\mathcal{V}$ representing the original problem $P_{0}$. The incumbent solution $\bar{x}\in\mathbb{Z}^{|\mathcal{I}|}\times\mathbb{R}^{n-|\mathcal{I}|}$ denotes the best feasible solution found at current iteration, its associated value $GU\!B=c^{\top}\bar{x}$ is called the global upper bound on the optimal value. The overall state of the optimization process is thus captured by the triplet $s=(\mathcal{V},\mathcal{E},\bar{x})$, we note $\mathcal{S}$ the set of all such triplets. [^1] To account for early resolution steps where no incumbent solution has yet been found, we define a special value for $\bar{x}$, whose $GU\!B=\infty$. For the sake of simplicity, we make this implicit in the remainder of the paper. Throughout the optimization process, B&B nodes are explored sequentially. We note $\mathcal{C}$ the set of visited or closed nodes, and $\mathcal{O}$ the set of unvisited or open nodes, such that $\mathcal{V}=\mathcal{C}\cup\mathcal{O}$. Figure 2 illustrates how B&B operates on an example. Originally, $\mathcal{O}=\{v_{0}\}$ and $\mathcal{C}=\emptyset$. At each iteration $t\geq 0$, the node selection policy $\rho:\mathcal{S}\rightarrow\mathcal{O}$ selects the next node to explore. Let $x^{*}_{LP}\in\mathbb{R}^{n}$ be the optimal solution to the linear program (LP) relaxation of $P_{t}$, the MILP associated with the current node $v_{t}$:
@@
Model-based reinforcement learning approaches exploit reversible access to the MDP to design policy improvement operators via planning (silver_general_2018; hansen2023td). Monte Carlo Tree Search (MCTS) is among the most widely used planning algorithms in MBRL. Given access to an environment model $(\mathcal{T},\mathcal{R})$, MCTS leverages value and policy estimates to select actions leading to promising states, while balancing an exploration-exploitation criterion. In order to extend the applicability of MCTS-based RL algorithms to broader control tasks where efficient simulators are not available, schrittwieser_mastering_2020 introduced MuZero, building upon the previous AlphaZero framework (silver_general_2018). Concretely, MuZero learns a model consisting of three interconnected networks. First, the representation network maps raw state observations $s_{t}$ to a latent state $\hat{s}_{t}=h(s_{t})$. This internal state can then be passed to the prediction network $f$ to obtain state policy and value estimates $(\hat{\mathrm{p}}_{t},\hat{\mathrm{v}}_{t})$. Alternatively, latent states can be passed along with an action $a_{t}$ to the dynamics network $g$, to produce predictions for both the true reward $r_{t}$ and the internal representation $h(s_{t+1})$ of the next visited state when taking action $a_{t}$ at $s_{t}$: $g(\hat{s}_{t},a_{t})=(\hat{r}_{t},\hat{s}_{t+1})$. At each decision step, MuZero integrates $h$, $f$ and $g$ to perform MCTS, thereby generating improved policy targets $\pi_{t}$ from which actions are sampled. During training, the model is unrolled for $K$ hypothetical steps and its predictions are aligned with sequences sampled from real trajectories collected by MCTS actors. By enforcing consistency between the predictions observed along simulated rollouts and those observed on real trajectories, MuZero constrains its model to capture only the information most relevant for accurately estimating future states’ policy and value. This approach was shown to significantly reduce the burden of MDP dynamics approximation, unlocking the use of MBRL in visually complex domains where model-free RL methods had previously dominated. Subsequent work introduced several enhancements to the MuZero framework, including a temporal consistency loss for the dynamics network (ye_mastering_2021) and the use of Gumbel search for improved search efficiency in MCTS (danihelka2022policy), which proved particularly useful in environments with large action space.
@@ 4 Related work
Following the seminal work by gasse_exact_2019, several contributions have proposed to build more complex neural network architectures based on transformers (lin_learning_2022) and recurrence mechanisms (seyfi_exact_2023) to improve the performance of IL branching agents, with limited success. In parallel, theoretical and computational analysis (bestuzheva_scip_2021; sun_improving_2022) have shown that neural networks trained by imitation could not rival the tree size achieved by strong branching (SB), the branching expert used in gasse_exact_2019. In fact, low tree sizes associated with SB turn out to be primarily due to the formulation improvements resulting from the massive number of LPs solved in SB, not to the intrinsic quality of the branching decisions themselves.
Since branching decisions are made sequentially, reinforcement learning appears as a natural candidate to learn good branching policies. Building on the works of etheve_reinforcement_2020 and scavuzzo_learning_2022, parsonson_reinforcement_2022 explored training RL branching agents on B&B trees expanded following solvers’ default node selection policies, rather than DFS. To address the partial observability induced by moving away from DFS, they trained their agent on retrospective trajectories, tree-diving trajectories extracted from original B&B episodes. More generally, a large body of work has proposed to learn, either by imitation or reinforcement, better-performing B&B heuristics outside of variable selection (nair_solving_2021; paulus_learning_2022; sonnerat_learning_2022). Building on the MDP formulation first proposed by he_learning_2014, recent RL contributions to primal search (sonnerat_learning_2022; wu_deep_2023), node selection (etheve_solving_2021; zhang_learning_2024), and cut selection (tang_reinforcement_2020; wang_learning_2023) have primarily focused on deploying traditional model-free RL algorithms within the B&B framework, adapting the action space to suit the specific heuristic being targeted.
Finally, machine learning applications to CO are not limited to B&B. For example, in the context of routing or scheduling problems, where exact resolution rapidly becomes prohibitive, agents are trained to learn direct search heuristics yielding high-quality feasible solutions (kool_attention_2019; chalumeau2023combinatorial). Drawing inspiration from model-based planning, pirnay2024self proposed a self-improvement operator that refines learning agents’ policies by generating alternative trajectories via stochastic beam search, and treating improved trajectories as targets for imitation learning. Applied to routing problems, their approach matched the performance of imitation learning agents trained on expert demonstrations, highlighting the potential of planning-based methods as a data-efficient alternative to expert supervision in combinatorial optimization.
@@ 5 Experimental study
We now assess the efficiency of our model-based branching agent, as we aim to answer the following questions.
@@
Set Covering Comb. Auction Max. Ind. Set Mult. Knapsack Norm. Score
Method Node Time Node Time Node Time Node Time Node Time
Presolve - $4.74$ - $0.90$ - $1.78$ - $0.20$ - -
Random $3289$ $5.94$ $1111$ $2.16$ $386.8$ $2.01$ $733.5$ $0.55$ $1068$ $454$
SB $35.8$ $12.93$ $28.2$ $6.21$ $24.9$ $45.87$ $161.7$ $0.69$ $46$ $4279$
SCIP $62.0$ $2.27$ $20.2$ $1.77$ $19.5$ $2.44$ $289.5$ $0.53$ $58$ $\color[rgb]{0,0,0}363$}
$\text{IL}^{\star}$} $\mathbf{133.8}$ $0.90$ $\mathbf{83.6}$ $0.65$ $\mathbf{40.1}$ $0.36$ $272.0$ $0.69$ $\mathbf{96}$ $116$
IL-DFS $136.4$ $\mathbf{0.74}$ $92.1$ $0.56$ $68.5$ $0.44$ $411.5$ $1.07$ $\color[rgb]{0,0,0}130$} $131$
PG-tMDP $649.4$ $2.32$ $168.0$ $0.94$ $153.6$ $0.92$ $436.9$ $1.57$ $272$ $254$
DQN-tMDP $\mathbf{175.8}$ $\mathbf{0.83}$ $203.3$ $1.11$ $168.0$ $1.00$ $266.4$ $0.73$ $207$ $188$
DQN-Retro $183.0$ $1.14$ $103.2$ $0.78$ $223.0$ $1.81$ $250.3$ $0.67$ $208$ $241$
PlanB&B 186.2 0.87 $\mathbf{84.7}$ $\mathbf{0.54}$ $\mathbf{44.8}$ $\mathbf{0.32}$ $\mathbf{220.0}$ $\mathbf{0.55}$ $\mathbf{100}$ $\mathbf{100}$
@@
Set Covering Comb. Auction Max. Ind. Set Mult. Knapsack Norm. Score
Method Node Time Node Time Node Time Node Time Node Time
Presolve - $\color[rgb]{0,0,0}12.3$} - $\color[rgb]{0,0,0}2.67$} - $\color[rgb]{0,0,0}5.16$} - $\color[rgb]{0,0,0}0.46$} - -
Random $271632$ $842$ $317235$ $749$ $215879$ $2102$ $93452$ $70.6$ $8050$ $3870$
SB $672.1$ $398$ $389.6$ $255$ $169.9$ $2172$ $\mathbf{1709}$ $\mathbf{12.5}$ $13$ $2243$
SCIP $3309$ $48.4$ $1376$ $14.77$ $3368$ $90.0$ $30620$ $22.1$ $121$ $132$
$\text{IL}^{\star}$} $\mathbf{2610}$ $23.1$ $\mathbf{1282}$ 9.4 $\mathbf{1993.0}$ $\mathbf{38.6}$ $11730$ $43.5$ $\mathbf{70}$ $\mathbf{83}$
IL-DFS $3103$ $\mathbf{22.5}$ $1828$ $10.2$ $3348$ $51.9$ $43705$ $130.8$ $151$ $136$
PG-tMDP $44649$ $221$ $6001$ $30.7$ $3133$ $43.6$ $35614$ $123$ $373$ $290$
DQN-tMDP $8632$ $71.3$ $20553$ $116$ $45634$ $477$ $22631$ $65.1$ $787$ $679$
DQN-Retro $6100$ $59.4$ $2908$ $18.4$ $119478$ $1863$ $27077$ $79.5$ $1166$ $1254$
PlanB&B $\mathbf{5869}$ $\mathbf{46.2}$ $\mathbf{1665}$ $\mathbf{9.1}$ $\mathbf{2853}$ $\mathbf{41.1}$ $\mathbf{13574}$ $\mathbf{51.2}$ $\mathbf{100}$ $\mathbf{100}$
@@
Table 1: Performance comparison of branching agents on four standard MILP benchmarks. For each method, we report total number of B&B nodes, presolve time and total solving time outside of presolve. Lower is better, red indicates best agent overall, blue indicates best among RL agents. Presolve is common to all methods. Following prior works, we report geometrical mean over 100 test instances unseen during training and over 100 higher-dimensional transfer instances. Norm. Score denotes the aggregate average performance obtained by each agent across the four MILP benchmarks, normalized by the score of PlanB&B.
@@
We consider four standard MILP benchmarks: set covering (SC), combinatorial auctions (CA), maximum independent set (MIS) and multiple knapsack (MK) problems. We use the open-source solver SCIP 8.0.3 (bestuzheva_scip_2021) as backend MILP solver, along with the Ecole library (prouvost_ecole_2020) for both instance generation and environment simulation. We train and test on instances of same dimensions as scavuzzo_learning_2022 and parsonson_reinforcement_2022, as described in Appendix A.
@@
The overall PlanB&B training pipeline is provided in Appendix F. Branching agents are trained on instances of each benchmark separately. For evaluation, we report performance in terms of both node count and solving time on 100 test instances unseen during training, as well as on 100 transfer instances of higher dimensions. All evaluation metrics are averaged over 5 random seeds. Importantly, when comparing ML-based branching strategies to standard SCIP heuristics such as RPB or SB, solving time remains the only reliable performance indicator. This is because invoking a custom branching rule in SCIP triggers auxiliary routines, such as conflict analysis and bound tightening, that strengthen the MILP formulation. While these mechanisms may reduce the number of explored nodes, they also incur computational overhead, often offsetting the gains in node efficiency and potentially increasing total solving time. As a result, node count becomes an unreliable proxy for performance, a limitation previously highlighted by gamrath2018measuring and scavuzzo_learning_2022.
@@
Computational results obtained on the four benchmarks are presented in Table 1. Standard deviations as well as further performance metrics are provided in Appendix I. Unlike conventional MBRL settings, PlanB&B operates under strict time and computational constraints. Crucially, every second spent on planning directly increases the overall solving time, thereby impacting final performance. Therefore, to allow systematic comparison, the results reported for PlanB&B in Table 1 reflect the performance of its policy network alone, without any computational budget allocated to MCTS simulations at evaluation time. On aggregate test instances, compared to prior RL baselines, PlanB&B’s policy network achieves $2\times$ reductions both in tree size and solving time. This performance gap broadens further to $3\times$ on aggregate transfer instances, underscoring the superior generalization capabilities of PlanB&B relative to prior RL agents. PlanB&B also outperforms the IL-DFS agent on both test and transfer instances, providing, to our knowledge, the first evidence of an RL-based branching strategy surpassing an IL agent trained to mimic strong branching. Specifically, our experiments demonstrate that, under a DFS node selection policy, PlanB&B learns branching strategies superior to that of the IL agent from gasse_exact_2019. When compared against the standard $\text{IL}^{\star}$} baseline, PlanB&B manages to achieve $\approx 10\%$ lower solving time on test instances, despite producing $\approx 5\%$ larger trees in average. However, on transfer instances, PlanB&B is globally unable to overcome the performance limitations imposed by DFS, although it still manages to outperform the $\text{IL}^{\star}$} baseline on combinatorial auction. The influence of DFS on the performance of branching agents is further analyzed in (Q4). Finally, PlanB&B clearly outperforms the default SCIP baseline on both test and transfer instances, despite operating under DFS.
@@
Given the performance trends observed in Figure 4, a natural question arises: does PlanB&B manage to derive better branching decisions than IL baselines by following SB more closely, or does it discover genuinely novel strategies? After all, as discussed in Appendix D, strong branching can be interpreted as a one-step MCTS procedure aimed at maximizing immediate dual gap reduction. To answer this, Table 2 reports several alignment metrics designed to assess which of the $\text{IL}^{\star}$} and PlanB&B policies more closely resembles strong branching. Across all metrics, PlanB&B policies exhibit lower alignment with SB than $\text{IL}^{\star}$} policies. Remarkably, the refined branching policy returned by the MCTS is roughly as close to SB as the policy yielded by the policy network. This suggests that PlanB&B surpasses IL baselines not through closer imitation of strong branching, but by discovering and exploiting original strategies.
SB ($\downarrow$) SB ($\uparrow$) SB ($\uparrow$)
@@
Table 3: $\text{IL}^{\star}$} agent discovery times, with $t_{r}=t_{d}/T$ on SC and MIS benchmarks.

### v2

@@ 2 Problem statement
with $n$ the number of variables, $m$ the number of linear constraints, $l,u\in\mathbb{R}^{n}$ the lower and upper bound vectors, $A\in\mathbb{R}^{m\times n}$ the constraint matrix, $b\in\mathbb{R}^{m}$ the right-hand side vector, $c\in\mathbb{R}^{n}$ the objective function, and $\mathcal{I}$ the indices of integer variables. In this work, we are interested in repeated MILPs of fixed dimension $\{P_{i}=(A_{i},b_{i},c_{i},l_{i},u_{i})\}_{i\in\mathcal{D}}$ sampled according to an unknown distribution $p_{0}$.
In order to solve MILPs efficiently, the B&B algorithm iteratively builds a binary tree $(\mathcal{V},\mathcal{E})$ where each node corresponds to a MILP, starting from the root node $v_{0}\in\mathcal{V}$ representing the original problem $P_{0}$. The incumbent solution $\bar{x}\in\mathbb{Z}^{|\mathcal{I}|}\times\mathbb{R}^{n-|\mathcal{I}|}$ denotes the best feasible solution found at current iteration, its associated value $GU\!B=c^{\top}\bar{x}$ is called the global upper bound on the optimal value. The overall state of the optimization process is thus captured by the triplet $s=(\mathcal{V},\mathcal{E},\bar{x})$, we note $\mathcal{S}$ the set of all such triplets. Throughout the optimization process, B&B nodes are explored sequentially. We note $\mathcal{C}$ the set of visited or closed nodes, and $\mathcal{O}$ the set of unvisited or open nodes, such that $\mathcal{V}=\mathcal{C}\cup\mathcal{O}$. Figure 2 illustrates how B&B operates on an example. Initially, $GU\!B=\infty$, $\mathcal{O}=\{v_{0}\}$, and $\mathcal{C}=\emptyset$. At each iteration $t\geq 0$, the node selection policy $\rho:\mathcal{S}\rightarrow\mathcal{O}$ selects the next node to explore. Let $x^{*}_{LP}\in\mathbb{R}^{n}$ be the optimal solution to the linear program (LP) relaxation of $P_{t}$, the MILP associated with the current node $v_{t}$:
@@
Model-based reinforcement learning approaches exploit reversible access to the MDP to design policy improvement operators via planning (silver_general_2018; hansen2023td). Monte Carlo Tree Search (MCTS) is among the most widely used planning algorithms in MBRL. Given access to an environment model $(\mathcal{T},\mathcal{R})$, MCTS leverages value and policy estimates to select actions leading to promising states, while balancing an exploration-exploitation criterion. In order to extend the applicability of MCTS-based RL algorithms to broader control tasks where efficient simulators are not available, schrittwieser_mastering_2020 introduced MuZero, building upon the previous AlphaZero framework (silver_general_2018). Concretely, MuZero learns a model consisting of three interconnected networks. First, the representation network $h$ maps raw state observations $s_{t}$ to a latent state $\hat{s}_{t}=h(s_{t})$. This internal state can then be passed to the prediction network $f$ to obtain state policy and value estimates $(\hat{\mathrm{p}}_{t},\hat{\mathrm{v}}_{t})$. Alternatively, latent states can be passed along with an action $a_{t}$ to the dynamics network $g$, to produce predictions for both the true reward $r_{t}$ and the internal representation $h(s_{t+1})$ of the next visited state when taking action $a_{t}$ at $s_{t}$: $g(\hat{s}_{t},a_{t})=(\hat{r}_{t},\hat{s}_{t+1})$. At each decision step, MuZero integrates $h$, $f$ and $g$ to perform MCTS, thereby generating improved policy targets $\pi_{t}$ from which actions are sampled. During training, the model is unrolled for $K$ hypothetical steps and its predictions are aligned with sequences sampled from real trajectories collected by MCTS actors. By enforcing consistency between the predictions observed along simulated rollouts and those observed on real trajectories, MuZero constrains its model to capture only the information most relevant for accurately estimating future states’ policy and value. This approach was shown to significantly reduce the burden of MDP dynamics approximation, unlocking the use of MBRL in visually complex domains where model-free RL methods had previously dominated. Subsequent work introduced several enhancements to the MuZero framework, including a temporal consistency loss for the dynamics network (ye_mastering_2021) and the use of Gumbel search for improved search efficiency in MCTS (danihelka2022policy), which proved particularly useful in environments with large action space.
@@ 4 Related work
Following the work of gasse_exact_2019, several studies have explored adopting more expressive neural architectures, from transformer-based models (lin_learning_2022) to recurrent designs (seyfi_exact_2023), for the purpose of learning to branch. However, these architectural extensions have yielded only limited gains. In parallel, theoretical and empirical analyses (dey_theoretical_2021) have shown that the small trees produced by strong branching (SB), the branching expert used in gasse_exact_2019, arise mainly from the extensive formulation tightening induced by solving large numbers of LPs, rather than from the intrinsic quality of its branching decisions.
Such findings cast doubt on the long-term effectiveness of purely imitation-based approaches. Since branching decisions unfold sequentially, reinforcement learning provides a principled alternative framework for learning to branch. Building on the works of etheve_reinforcement_2020 and scavuzzo_learning_2022, parsonson_reinforcement_2022 explored training RL branching agents on B&B trees expanded following solvers’ default node selection policies, rather than DFS. To address the partial observability induced by moving away from DFS, they trained their agent on retrospective trajectories, tree-diving trajectories extracted from original B&B episodes.
More generally, a substantial line of research (nair_solving_2021; paulus_learning_2022) has explored augmenting B&B heuristics with machine learning algorithms beyond variable selection. Building on the MDP formulation first proposed by he_learning_2014, recent RL contributions to primal search (sonnerat_learning_2022; wu_deep_2023), node selection (etheve_solving_2021; zhang_learning_2024), and cut selection (tang_reinforcement_2020; wang_learning_2023) have primarily focused on deploying traditional model-free RL algorithms within the B&B framework, adapting the action space to suit the specific heuristic being targeted.
Finally, machine learning methods have been applied beyond the scope of exact combinatorial optimization. In large routing and scheduling problems, where exact resolution quickly becomes intractable, recent works have trained neural agents to act as direct-search heuristics, capable of producing strong feasible solutions (kool_attention_2019; chalumeau2023combinatorial). Drawing inspiration from model-based planning, pirnay2024self proposed a self-improvement operator that refines learning agents’ policies by generating alternative trajectories via stochastic beam search, and treating improved trajectories as targets for imitation learning. Applied to routing problems, their approach matched the performance of imitation learning agents trained on expert demonstrations, highlighting the potential of planning-based methods as a data-efficient alternative to expert supervision in combinatorial optimization.
@@ 5 Experimental study
We now assess the efficiency of our model-based branching agent, as we aim to answer the following questions:
@@
Set Covering Comb. Auction Max. Ind. Set Mult. Knapsack Norm. Score $\mathbf{\pm}$ Conf. Interval
Method Node Time Node Time Node Time Node Time Node CI Time CI
Presolve - $4.74$ - $0.90$ - $1.78$ - $0.20$ - - - -
Random $3289$ $5.94$ $1111$ $2.16$ $386.8$ $2.01$ $733.5$ $0.55$ $1068$ $\pm 113$ $454$ $\pm 51$
SB $35.8$ $12.93$ $28.2$ $6.21$ $24.9$ $45.87$ $161.7$ $0.69$ $46$ $\pm 0$ $4279$ $\pm 57$
SCIP $62.0$ $2.27$ $20.2$ $1.77$ $19.5$ $2.44$ $289.5$ $0.53$ $58$ $\pm 0$ $\color[rgb]{0,0,0}363$} $\pm 3$
$\text{IL}^{\star}$} $\mathbf{133.8}$ $0.90$ $\mathbf{83.6}$ $0.65$ $\mathbf{40.1}$ $0.36$ $272.0$ $0.69$ $\mathbf{96}$ $\mathbf{\pm 17}$ $116$ $\pm 15$
IL-DFS $136.4$ $\mathbf{0.74}$ $92.1$ $0.56$ $68.5$ $0.44$ $411.5$ $1.07$ $\color[rgb]{0,0,0}130$} $\pm 26$ $131$ $\pm 20$
PG-tMDP $649.4$ $2.32$ $168.0$ $0.94$ $153.6$ $0.92$ $436.9$ $1.57$ $272$ $\pm 46$ $254$ $\pm 50$
DQN-tMDP $\mathbf{175.8}$ $\mathbf{0.83}$ $203.3$ $1.11$ $168.0$ $1.00$ $266.4$ $0.73$ $207$ $\pm 25$ $188$ $\pm 16$
DQN-Retro $183.0$ $1.14$ $103.2$ $0.78$ $223.0$ $1.81$ $250.3$ $0.67$ $208$ $\pm 23$ $241$ $\pm 25$
PlanB&B 186.2 0.87 $\mathbf{84.7}$ $\mathbf{0.54}$ $\mathbf{44.8}$ $\mathbf{0.32}$ $\mathbf{220.0}$ $\mathbf{0.55}$ $\mathbf{100}$ $\mathbf{\pm 9}$ $\mathbf{100}$ $\mathbf{\pm 12}$
@@
Set Covering Comb. Auction Max. Ind. Set Mult. Knapsack Norm. Score $\mathbf{\pm}$ Conf. Interval
Method Node Time Node Time Node Time Node Time Node CI Time CI
Presolve - $\color[rgb]{0,0,0}12.3$} - $\color[rgb]{0,0,0}2.67$} - $\color[rgb]{0,0,0}5.16$} - $\color[rgb]{0,0,0}0.46$} - - - -
Random $271632$ $842$ $317235$ $749$ $215879$ $2102$ $93452$ $70.6$ $8050$ $\pm 1646$ $3870$ $\pm 847$
SB $672.1$ $398$ $389.6$ $255$ $169.9$ $2172$ $\mathbf{1709}$ $\mathbf{12.5}$ $13$ $\pm 0$ $2243$ $\pm 47$
SCIP $3309$ $48.4$ $1376$ $14.77$ $3368$ $90.0$ $30620$ $22.1$ $121$ $\pm 0$ $132$ $\pm 1$
$\text{IL}^{\star}$} $\mathbf{2610}$ $23.1$ $\mathbf{1282}$ 9.4 $\mathbf{1993.0}$ $\mathbf{38.6}$ $11730$ $43.5$ $\mathbf{70}$ $\mathbf{\pm 7}$ $\mathbf{83}$ $\mathbf{\pm 7}$
IL-DFS $3103$ $\mathbf{22.5}$ $1828$ $10.2$ $3348$ $51.9$ $43705$ $130.8$ $151$ $\pm 30$ $136$ $\pm 21$
PG-tMDP $44649$ $221$ $6001$ $30.7$ $3133$ $43.6$ $35614$ $123$ $373$ $\pm 48$ $290$ $\pm 42$
DQN-tMDP $8632$ $71.3$ $20553$ $116$ $45634$ $477$ $22631$ $65.1$ $787$ $\pm 128$ $679$ $\pm 78$
DQN-Retro $6100$ $59.4$ $2908$ $18.4$ $119478$ $1863$ $27077$ $79.5$ $1166$ $\pm 256$ $1254$ $\pm 218$
PlanB&B $\mathbf{5869}$ $\mathbf{46.2}$ $\mathbf{1665}$ $\mathbf{9.1}$ $\mathbf{2853}$ $\mathbf{41.1}$ $\mathbf{13574}$ $\mathbf{51.2}$ $\mathbf{100}$ $\mathbf{\pm 9}$ $\mathbf{100}$ $\mathbf{\pm 8}$
@@
Table 1: Performance comparison of branching agents on four standard MILP benchmarks. For each method, we report the total number of B&B nodes, presolve time, and total solving time excluding presolve. The presolve phase is identical across all methods. Lower values indicate better performance. Red highlights the overall best agent, while blue marks the best-performing RL-based agent. Following prior work, results are reported as the geometric mean over 100 unseen test instances and an additional 100 higher-dimensional transfer instances, averaged across 5 random seeds. Norm. Score represents the aggregate average performance of each agent across the four MILP benchmarks, normalized by the score of PlanB&B. Aggregate confidence intervals (CI) are reported for each baselines; additional per-benchmark confidence intervals are provided in Appendix I.
@@
We consider four standard MILP benchmarks: set covering (SC), combinatorial auctions (CA), maximum independent set (MIS) and multiple knapsack (MK) problems. In our experiments, SCIP 8.0.3 (bestuzheva_scip_2021) is used as backend MILP solver, along with the Ecole library (prouvost_ecole_2020) for instance generation, see Appendix A for further benchmark details.
@@
The overall PlanB&B training pipeline is provided in Appendix F. Branching agents are trained on instances of each benchmark separately. For evaluation, we report performance in terms of both node count and solving time on 100 test instances unseen during training, as well as on 100 transfer instances of higher dimensions. Evaluation metrics are averaged over 5 random seeds. Importantly, when comparing ML-based branching strategies to standard SCIP heuristics such as RPB or SB, solving time remains the only reliable performance indicator. This is because invoking a custom branching rule in SCIP triggers auxiliary routines, such as conflict analysis and bound tightening, that strengthen the MILP formulation while incurring computational overhead.
@@
Computational results obtained on the four benchmarks are presented in Table 1. Standard deviations, additional performance metrics as well as targeted ablations are provided in Appendix I. Unlike conventional MBRL settings, PlanB&B operates under strict time and computational constraints. Crucially, every second spent on planning directly increases the overall solving time, thereby impacting final performance. Therefore, to allow systematic comparison, the results reported for PlanB&B in Table 1 reflect the performance of its policy network alone, without any computational budget allocated to MCTS simulations at evaluation time. On aggregate test instances, compared to prior RL baselines, PlanB&B’s policy network achieves $2\times$ reductions both in tree size and solving time. This performance gap broadens further to $3\times$ on aggregate transfer instances, underscoring the superior generalization capabilities of PlanB&B relative to prior RL agents. PlanB&B also outperforms the IL-DFS agent on both test and transfer instances, providing, to our knowledge, the first evidence of an RL-based branching strategy surpassing an IL agent trained to mimic strong branching. Specifically, our experiments demonstrate that, under a DFS node selection policy, PlanB&B learns branching strategies superior to that of the IL agent from gasse_exact_2019. When compared against the standard $\text{IL}^{\star}$} baseline, PlanB&B manages to achieve $\approx 10\%$ lower solving time on test instances, despite producing $\approx 5\%$ larger trees in average. However, on transfer instances, PlanB&B is globally unable to overcome the performance limitations imposed by DFS, although it still manages to outperform the $\text{IL}^{\star}$} baseline on combinatorial auction. The influence of DFS on the performance of branching agents is further analyzed in (Q4). Finally, PlanB&B clearly outperforms the default SCIP baseline on both test and transfer instances, despite operating under DFS, a performance achieved by no other learning-based baseline.
@@
Given the performance trends observed in Figure 4, a natural question arises: does PlanB&B manage to derive better branching decisions than IL baselines by following SB more closely, or does it discover genuinely novel strategies? After all, strong branching can be interpreted as a one-step MCTS procedure aimed at maximizing immediate dual gap reduction. To answer this, Table 2 reports several alignment metrics designed to assess which of the $\text{IL}^{\star}$} and PlanB&B policies more closely resembles strong branching. Across all metrics, PlanB&B policies exhibit lower alignment with SB than $\text{IL}^{\star}$} policies. Remarkably, the refined branching policy returned by the MCTS is roughly as close to SB as the policy yielded by the policy network. This suggests that PlanB&B surpasses IL baselines not through closer imitation of strong branching, but by discovering and exploiting original strategies.
($\downarrow$) SB ($\uparrow$) SB ($\uparrow$) SB
@@
Table 3: $\text{IL}^{\star}$} agent discovery times, with $t_{r}=t_{d}/T$ on SC and MIS benchmarks, see Appendix I for further results.